# Memory effects in a dynamical decoupling process by parity kicks

### v1

@@ III.2 Memory effects with specific time instants
Intuitively, there exists some special time instants in a dynamical decoupling process that are closely related to the environmental memory effects. For example, in one cycle of parity kicks described by Eq. (7), the information about the system’s history in $[0,\tau]$ [stored in $\rho_{SE}(\tau)$] is exploited to time-reverse the dynamics after the first kick and ultimately eliminate the influence of environment at $2\tau$ (after the second kick). This motivates us to focus on the memory effects associated with specific time instants (e.g., $t_{1}=\tau$ and $t_{2}=2\tau$ in a parity kick cycle), rather than maximize the memory effects over all the possible $t_{1}$ and $t_{2}$ in a time interval like $[0,\infty]$ or $[0,t_{max}]$. Such a treatment is expected to help us establish simple and quantitative connections between the dynamical decoupling characteristics and strength of memory effects. Moreover, the computational effort would be significantly reduced since the maximization over the time instants are not required.
@@
To evaluate the strength of memory effects in a dynamical decoupling process with more than 2 kicks, e.g., more than one cycle of Eq. (7), we extend Eq. (26) and (27) to their multi-time forms. Notice that the divisibility condition Eq. (19) for a quantum Markov process can be extended to
$\!\!\!\!\!\!\!\!\!\!\Lambda_{M}(t_{n},t_{0})=\Lambda_{M}(t_{n},t_{n\!-\!1})\cdots\Lambda_{M}(t_{2},t_{1})\Lambda_{M}(t_{1},t_{0})$ (28)
for $t_{0}\leqslant t_{1}\cdots\leqslant t_{n}$. Following similar procedures discussed above, we arrive at the non-Markovian multi-time inequality in terms of the dynamical map $T(t_{b},t_{a})$:
@@
The multi-time inequality can also be interpreted as memory effects from the perspective of the environment as follows. Let the left-hand and the right-hand sides of Eq. (29) act on a system initial state $\rho_{S}(t_{0})$ so that $\rho_{S}(t_{n})=T(t_{n},t_{0})\rho_{S}(t_{0})$ and $\rho^{\prime}_{S}(t_{2})=T(t_{n},t_{n-1})T(t_{n-1},t_{n-2})\cdots T(t_{1},t_{0})\rho_{S}(t_{1})$. The system-environment evolutions described by the latter equation can be understood as that by the former equation subjecting to $n-1$ initializations of the environment state at $t_{1,2,\cdots,n-1}$. The initialization at $t_{j}$ erases the system’s history in the time interval $[t_{j-1},t_{j}]$ as done in Eq. (23). Therefore, $\rho_{S}(t_{n})\neq\rho^{\prime}_{S}(t_{n})$ signifies that the environment (as well as the system-environment correlations) remembers the system’s histories which in turn influences the future system at $t_{n}$. Then, the strength of memory effects with $n$ specific time instants $t_{1,2,\cdots,n}$ ($t_{0}$ is fixed) is given by
@@
$N_{M}^{t_{1:n}}[\rho_{S}(t_{0})]=D[\rho_{S}(t_{2}),\rho^{\prime}_{S}(t_{2})]$ (31)
where $\rho_{S}(t_{2})$ and $\rho^{\prime}_{S}(t_{2})$ are obtained from Eq. (29) as mentioned before. Then the strengths of memory effects with many specific time instants can be applied to a general dynamical decoupling processes described by one or more cycles of Eq. (9) by letting $t_{0}=0$, and $t_{j}=j\tau$ (the time instant after each kick). Within the definitions Eq. (30) and (31), it is easy to see that $0\leqslant N_{M}^{t_{1:n}}\leqslant 1$ and $0\leqslant N_{M}^{t_{1:n}}[\rho_{S}(t_{0})]\leqslant 1$ are satisfied. However, the values of $N_{M}^{t_{1:2}}$ and $N_{M}^{t_{1:n}}$ may not be compared directly since the number of environment initializations are different in the two measures, so do $N_{M}^{t_{1:2}}[\rho_{S}(t_{0})]$ and $N_{M}^{t_{1:n}}[\rho_{S}(t_{0})]$.
@@ IV Memory effects and dynamical decoupling
It is known that a non-Markovian environment is necessary for the success of a dynamical decoupling control of an open quantum system. Therefore, a system with weak memory effects may lead to failure of dynamical decoupling. On the other hand, in an ideal dynamical decoupling process where infinitely frequent ($t_{c}\rightarrow 0$) instantaneous kicks act on the system, the system may be perfectly decoupled from the environment so that the controlled dynamics becomes effectively unitary and thus Markovian. These seemingly contradictive results motivate us to consider the memory effects in both uncontrolled and controlled processes to better understand their roles in a dynamical decoupling process.
@@
$=$ $\text{Tr}_{E}[U(t_{b},t_{a})\rho_{S}(t_{a})\otimes\rho_{E}(t_{a})U(t_{b},t_{a})^{{\dagger}}]$
[…]
@@
as the effect of a dynamical decoupling control (from 00 to $t_{f}=kt_{c}$) on the dynamical map $T(t_{f},0)$. It represents the difference between the controlled (practically attainable) dynamical map $\tilde{T}(t_{f},0)$ and the uncontrolled one $\bar{T}(t_{f},0)$. Correspondingly, the effect of a dynamical decoupling control on the final state is given by
@@ IV.1 One cycle of parity kicks
Let $\tilde{N}_{M}^{t_{1:2}}$ denote the strength of memory effects for one cycle of parity kicks where $t_{1}=\tau$ and $t_{2}=t_{c}=2\tau$ (after each kick) and $t_{0}=0$. Using the condition Eq. (34), $\tilde{N}_{M}^{t_{1:2}}$ satisfies
$\tilde{N}_{M}^{t_{1:2}}$ $=$ $D[\rho_{\tilde{T}(2\tau,0)},\rho_{\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}]$ (37)
$=$ $D[\rho_{\tilde{T}(2\tau,0)},\rho_{\mathcal{K}_{P}\bar{T}(2\tau,\tau)\mathcal{K}_{P}\bar{T}(\tau,0)}].$
@@
It is observed that if $\mathcal{K}_{P}$ commutes with $\bar{T}(t_{b},t_{a})$, i.e., $\mathcal{K}_{P}\bar{T}(t_{b},t_{a})=\bar{T}(t_{b},t_{a})\mathcal{K}_{P}$, then the effect of two parity kicks $E=D[\rho_{\tilde{T}(2\tau,0)},\rho_{\bar{T}(2\tau,0)}]$ is bounded by $\tilde{N}_{M}^{t_{1:2}}$ and $\bar{N}_{M}^{t_{1:2}}$ as
@@
which is a main finding of our work. The relation can be straightforwardly proved by virtue of the condition $\mathcal{K}_{P}T(2\tau,\tau)\mathcal{K}_{P}T(\tau,0)=\mathcal{K}_{P}\mathcal{K}_{P}\bar{T}(2\tau,\tau)\bar{T}(\tau,0)=\bar{T}(2\tau,\tau)\bar{T}(\tau,0)$ and the triangle inequality
@@
that the trace distance obeys. The above result shows that under certain conditions, the upper and lower bounds of the control effect (on the final dynamical map) is determined by the strength of memory effects in both the controlled and uncontrolled process. Meanwhile, the difference between the controlled and uncontrolled dynamical map is a manifestation of memory effects since $E=0$ if $\tilde{N}_{M}^{t_{1:2}}=\bar{N}_{M}^{t_{1:2}}=0$.
If we are interested in the quantum states rather than the dynamical maps in a cycle of parity kicks, the above results can be formulated following similar steps. As discussed above, the strength of memory effects in one cycle of parity kicks with the initial state $\rho_{S}(0)$ is given by
@@
where $\tilde{\rho}_{S}(t_{c})=\tilde{T}(2\tau,0)\rho_{S}(0)$ and $\tilde{\rho}^{\prime}_{S}(t_{c})=\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)\rho_{S}(0)=\mathcal{K}_{P}\bar{T}(2\tau,\tau)\mathcal{K}_{P}\bar{T}(\tau,0)\rho_{S}(0)$. In the absence of control, the strength of memory effects with the same time instants is
@@
if $\mathcal{K}_{P}$ commutes with $\bar{T}(t_{b},t_{a})$. Similarly, the change of the final states (brought by control) is upper and lower bounded by the memory effects in the controlled and uncontrolled process. Meanwhile, $E[\rho_{S}(0)]>0$ is a manifestation of memory effects.
@@ IV.2 General case
Our theory can be generated to a general dynamical decoupling process with $n$ kicks ($n\geqslant 2$), specifically, one or more cycles of Eq. (9) where the kicks are designed by Eq. (14). Consider a dynamical decoupling process (from 0 to $t_{f}$) with $k$ cycles while each cycle contains $n_{g}$ kicks. When $k>1$, the kicks satisfies $K_{j+n_{g}}=K_{j}$. The process can be described by $\tilde{T}(n\tau,0)$ where $n=kn_{g}$. Let $t_{0}=0$ and $t_{j}=j\tau$ (after each kick), then the strength of memory effects in the dynamical decoupling process satisfies
@@
$=$ $\!D\{\!\rho_{\tilde{T}(n\tau,0)},\rho_{\mathcal{K}_{n}\!\bar{T}[n\tau,(\!n\!-\!1\!)\tau]\cdots\mathcal{K}_{2}\!\bar{T}(2\tau,\tau)\mathcal{K}_{1}\!\bar{T}(\tau,0)}\!\}$
following the definition Eq. (30) and the condition Eq. (34). Using the same time instants $t_{1,2,\cdots,n}$, the strength of memory effects without control is given by
@@
With the condition Eq. (15) and the triangle inequality of trace distance, it is straightforward to verify that if each $\mathcal{K}_{j}$ commutes with $\bar{T}(t_{b},t_{a})$, then the effect of control on the dynamical map satisfies
$|\tilde{N}_{M}^{t_{1:n}}-\bar{N}_{M}^{t_{1:n}}|\leqslant E\leqslant\tilde{N}_{M}^{t_{1:n}}+\bar{N}_{M}^{t_{1:n}}.$ (46)
The result reduces to Eq.(39) when $n=2$ ($k=1$ and $n_{g}=2$) and can be applied to more than one cycle of parity kicks ($k>1$ and $n_{g}=2$). If $n>2$, the interpretation of $\tilde{N}_{M}^{t_{1:n}}$ ($\bar{N}_{M}^{t_{1:n}}$) as memory effects is slightly different with that of $\tilde{N}_{M}^{t_{1:2}}$ ($\bar{N}_{M}^{t_{1:2}}$) as discussed in Sec. III B.
@@
where $\tilde{\rho}_{S}(kt_{c})=\tilde{T}(n\tau,0)\rho_{S}(0)$ and $\tilde{\rho}^{\prime}_{S}(kt_{c})=\tilde{T}[n\tau,(n-1)\tau]\cdots\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)\rho_{S}(0)=\mathcal{K}_{n}\tilde{T}[n\tau,(n-1)\tau]\cdots\mathcal{K}_{2}T(2\tau,\tau)\mathcal{K}_{1}T(\tau,0)\rho_{S}(0)$. In the uncontrolled process, the strength of memory effects is
@@
if each $\mathcal{K}_{j}$ commutes with $\bar{T}(t_{b},t_{a})$.

### v2

@@ III.2 Memory effects with specific time instants
Intuitively, there exists some special time instants in a dynamical decoupling process that are closely related to the environmental memory effects. For example, in one cycle of parity kicks described by Eq. (7), the information about the system’s history in $0,\tau] [$stored in $\rho_{S\!E}(\tau)$] is exploited to time-reverse the dynamics after the first kick and ultimately eliminate the influence of environment at $2\tau$ (after the second kick). This motivates us to focus on the memory effects associated with specific time instants (e.g., $t_{1}=\tau$ and $t_{2}=2\tau$ in a parity kick cycle), rather than maximize the memory effects over all the possible $t_{1}$ and $t_{2}$ in a time interval like $0,\infty]$ as in Ref. [35] or a finite one $0,t_{max}]$ as in Ref. [36]. Such a treatment is expected to help us establish simple and quantitative connections between the dynamical decoupling characteristics and strength of memory effects. Moreover, the computational effort would be significantly reduced since the maximization over the time instants are not required.
@@
To evaluate the strength of memory effects in a dynamical decoupling process with more than 2 kicks, e.g., more than one cycle of Eq. (7), we extend Eq. (23) and (24) to their multi-time forms. In a Markovian quantum process, by repeatedly applying the divisibility Eq. (19) to one of the terms on the right-hand side of Eq. (19), there are
$\!\!\!\!\!\!\!\!\!\!T(t_{n},t_{0})=T(t_{n},t_{n\!-\!1})\cdots T(t_{2},t_{1})T(t_{1},t_{0})$ (25)
for $t_{0}\leqslant t_{1}\cdots\leqslant t_{n}$. Similarly, the violation of the multi-time divisibility condition is manifested by the inequality:
@@
The multi-time inequality can also be interpreted as memory effects from the perspective of the environment as follows. Let the left-hand and the right-hand sides of Eq. (26) act on a system initial state $\rho_{S}(t_{0})$ so that $\rho_{S}(t_{n})=T(t_{n},t_{0})\rho_{S}(t_{0})$ and $\rho^{\prime}_{S}(t_{n})=T(t_{n},t_{n-1})\cdots T(t_{2},t_{1})T(t_{1},t_{0})\rho_{S}(t_{0})$. The system-environment evolutions described by the latter equation can be understood as that by the former equation subjecting to n1n-1 initializations of the environment state at $t_{1,2,\cdots,n-1}$. The initialization at $t_{j}$ erases the system’s history in the time interval $t_{j-1},t_{j}]$ by the process $\rho_{S\!E}(t_{j})\rightarrow\rho_{S}(t_{j})\otimes\rho_{E}(t_{j})$ with $t_{j}=t_{I}$. Therefore, $\rho_{S}(t_{n})\neq\rho^{\prime}_{S}(t_{n})$ signifies that the environment (as well as the system-environment correlations) remembers the system’s histories which in turn influences the future system at $t_{n}$. Then, the strength of memory effects (in terms of dynamical maps) with $n$ specific time instants $t_{1,2,\cdots,n}$ ($t_{0}$ is fixed) is given by
@@
$N_{M}^{t_{1:n}}[\rho_{S}(t_{0})]=D[\rho_{S}(t_{n}),\rho^{\prime}_{S}(t_{n})]$ (28)
where $\rho_{S}(t_{n})$ and $\rho^{\prime}_{S}(t_{n})$ are obtained from Eq. (26) as mentioned before. Then the strengths of memory effects with many specific time instants can be applied to a general dynamical decoupling processes described by one or more cycles of Eq. (9) by letting $t_{0}=0$, and $t_{j}=j\tau$ (the time instant after each kick). Within the definitions Eq. (27) and (28), it is easy to see that $0\leqslant N_{M}^{t_{1:n}}\leqslant 1$ and $0\leqslant N_{M}^{t_{1:n}}[\rho_{S}(t_{0})]\leqslant 1$ are satisfied. However, the values of $N_{M}^{t_{1:2}}$ and $N_{M}^{t_{1:n}}$ may not be compared directly since the number of environment initializations are different in the two measures, so do $N_{M}^{t_{1:2}}[\rho_{S}(t_{0})]$ and $N_{M}^{t_{1:n}}[\rho_{S}(t_{0})]$.
@@ IV Memory effects and dynamical decoupling
It is known that a non-Markovian environment is necessary for the success of a dynamical decoupling control of an open quantum system. Therefore, a system with weak memory effects may lead to failure of dynamical decoupling. On the other hand, in an ideal dynamical decoupling process where infinitely frequent ($t_{c}\rightarrow 0$) instantaneous kicks act on the system, the system may be perfectly decoupled from the environment so that the controlled dynamics becomes effectively unitary and thus Markovian. Thus we may not find a monotonic relationship between the performance of dynamical decoupling and the strength of memory effects in the controlled dynamics. In the following, we consider the memory effects as well as performances in both uncontrolled and controlled process to investigate their quantitative relationship.
@@
$=$ $\mathrm{Tr}_{E}[U(t_{b},t_{a})\rho_{S}(t_{a})\otimes\rho_{E}(t_{a})U(t_{b},t_{a})^{{\dagger}}]$
where $U(t_{b},t_{a})=\mathcal{T}e^{-i\int_{t_{a}}^{t_{b}}H_{c}(t^{\prime})dt^{\prime}}$. For example, for a cycle of ideal parity kicks starting at $t_{a}=0$, there is $U(\tau,0)=U(2\tau,\tau)=K_{P}e^{-iH_{0}\tau}$.
@@
as the effect of a dynamical decoupling control (from 0 to $t_{f}=kt_{c}$) on the dynamical map $T(t_{f},0)$. It represents the difference between the controlled (practically attainable) dynamical map $\tilde{T}(t_{f},0)$ and the uncontrolled one $\bar{T}(t_{f},0)$. Correspondingly, the effect of a dynamical decoupling control on the final state is given by
@@ IV.1 One cycle of parity kicks
Let $\tilde{N}_{M}^{t_{1:2}}$ denote the strength of memory effects for one cycle of parity kicks where $t_{1}=\tau$ and $t_{2}=t_{c}=2\tau$ (after each kick) and $t_{0}=0$, i.e.,
$\tilde{N}_{M}^{t_{1:2}}$ $=$ $D[\rho_{\tilde{T}(2\tau,0)},\rho_{\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}].$ (34)
@@
It is observed that if $\mathcal{K}_{P}$ commutes with $\bar{T}(t_{b},t_{a})$, i.e.,
$\mathcal{K}_{P}\bar{T}(t_{b},t_{a})=\bar{T}(t_{b},t_{a})\mathcal{K}_{P},$ (36)
then the effect of two parity kicks on the dynamical map, i.e., $E=D[\rho_{\tilde{T}(2\tau,0)},\rho_{\bar{T}(2\tau,0)}]$, is bounded by $\tilde{N}_{M}^{t_{1:2}}$ and $\bar{N}_{M}^{t_{1:2}}$ as
@@
which is a main finding of our work. The condition Eq. (37) can be derived with the help of Eq. (31) and the triangle inequality
@@
that the trace distance obeys. The derivation of is presented in Appendix A. The above result shows that under certain conditions, the upper and lower bounds of the control effect (on the final dynamical map) is determined by the strength of memory effects in both the controlled and uncontrolled process. Meanwhile, the difference between the controlled and uncontrolled dynamical map is a manifestation of memory effects since $E=0$ if $\tilde{N}_{M}^{t_{1:2}}=\bar{N}_{M}^{t_{1:2}}=0$.
If we are interested in the quantum states rather than the dynamical maps in a cycle of parity kicks, the bounds of the memory effect strength can be obtained following similar steps. As discussed above, the strength of memory effects for one cycle of parity kicks with the initial state $\rho_{S}(0)$ is given by
@@
where $\tilde{\rho}_{S}(t_{c})=\tilde{T}(2\tau,0)\rho_{S}(0)$ and $\tilde{\rho}^{\prime}_{S}(t_{c})=\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)\rho_{S}(0)$. In the absence of control, the strength of memory effects with the same time instants is given by
@@
if $\mathcal{K}_{P}$ commutes with $\bar{T}(t_{b},t_{a})$ (see Appendix A for its derivation). Similarly, the change of the final states (brought by control) is upper and lower bounded by the memory effects in the controlled and uncontrolled process. Meanwhile, $E[\rho_{S}(0)]>0$ is a manifestation of memory effects.
@@ IV.2 General case
Our theory can be extended to a dynamical decoupling process with $n$ kicks ($n\geqslant 2$), specifically, one or more cycles of Eq. (9) where the kicks are designed by Eq. (14). Consider a dynamical decoupling process (from 0 to $t_{f}$) with $k$ cycles while each cycle contains $n_{g}$ kicks. When $k>1$, the kicks satisfies $K_{j+n_{g}}=K_{j}$. The process can be described by $\tilde{T}(n\tau,0)$ where $n=kn_{g}$. Let $t_{0}=0$ and $t_{j}=j\tau$ (after each kick), then the strength of memory effects in the dynamical decoupling process satisfies
@@
following the definition Eq. (27) Using the same time instants $t_{1,2,\cdots,n}$, the strength of memory effects without control is given by
@@
With the condition Eq. (15) and the inequality Eq. (38), it is found that if each $\mathcal{K}_{j}$ commutes with $\bar{T}(t_{b},t_{a})$, then the effect of control on the dynamical map satisfies
$|\tilde{N}_{M}^{t_{1:n}}-\bar{N}_{M}^{t_{1:n}}|\leqslant E\leqslant\tilde{N}_{M}^{t_{1:n}}+\bar{N}_{M}^{t_{1:n}}$ (44)
where $E=D[\rho_{\tilde{T}(n\tau,0)},\rho_{\bar{T}(n\tau,0)}]$. The derivation of Eq. (44) is presented in Appendix B. The result reduces to Eq. (37) when $n=2$ ($k=1$ and $n_{g}=2$) and can be applied to more than one cycle of parity kicks ($k>1$ and $n_{g}=2$). If $n>2$, the interpretation of $\tilde{N}_{M}^{t_{1:n}}$ ($\bar{N}_{M}^{t_{1:n}}$) as memory effects is slightly different with that of $\tilde{N}_{M}^{t_{1:2}}$ ($\bar{N}_{M}^{t_{1:2}}$) as discussed in Sec. III B.
@@
where $\tilde{\rho}_{S}(kt_{c})=\tilde{T}(n\tau,0)\rho_{S}(0)$ and $\tilde{\rho}^{\prime}_{S}(kt_{c})=\tilde{T}[n\tau,(n-1)\tau]\cdots\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)\rho_{S}(0)$. In the uncontrolled process, the strength of memory effects is
@@
if each $\mathcal{K}_{j}$ commutes with $\bar{T}(t_{b},t_{a})$ (see Appendix B for its derivation).
The general condition (44) ($n\geqslant 2$) can be experimentally verified by quantum process tomography [44] of the corresponding dynamical maps. If $\mathcal{K}_{j}$ commutes with $\bar{T}(t_{b},t_{a})$, there is $\tilde{T}[n\tau,(n\!-\!1)\tau]\cdots\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)=\bar{T}[n\tau,(n\!-\!1)\tau]\cdots\bar{T}(2\tau,\tau)\bar{T}(\tau,0)$ (see Appendix B). Then the characterizations of $\bar{T}(n\tau,0)$, $\bar{T}[n\tau,(n\!-\!1)\tau]\cdots\bar{T}(2\tau,\tau)\bar{T}(\tau,0)$ and $\tilde{T}(n\tau,0)$ are adequate to calculate the Choi-Jamiółkowski matrices in Eq. (44). Typically, this requires the tomography of $n+2$ dynamical maps. Alternatively, using $|\psi_{A\!S}\rangle$ (defined in Sec. III.1) as the initial state, $\rho_{\Lambda}$ can be obtained by quantum state tomography of the final state of the ASA\!-\!S composite system after $\mathbb{I}\otimes\Lambda$ [45]. Similarly, the condition (47) can be experimentally verified by quantum state tomography of the final states, $\bar{\rho}_{S}(kt_{c})$, $\bar{\rho}^{\prime}_{S}(kt_{c})$ and $\tilde{\rho}_{S}(kt_{c})$ while a particular initial state $\rho_{S}(0)$ is used. These final states can also be calculated by their corresponding dynamical maps obtained from quantum process tomography. In the above cases, if $\rho_{E}(t_{a})$ is stationary under $H_{E}$ (assuming $H_{0}$ is time-dependent), there is $\bar{T}(t_{b},t_{a})=\bar{T}(t_{b}-t_{a},0)$ [46]. Then the characterization of $\bar{T}[n\tau,(n\!-\!1)\tau]\cdots\bar{T}(2\tau,\tau)\bar{T}(\tau,0)$ only requires the tomography of $\bar{T}(\tau,0)$.
The existence of the commutation relation between $K_{j}$ and $\bar{T}(t_{b},t_{a})$ relies on the kick $K_{j}$ as well as the dynamical map $\bar{T}(t_{b},t_{a})$ which is a function of the system-environment Hamiltonian and the environment initial state. In general, the commutation relation can be verified with the analytical and numerical solution of the open dynamics, or by experiments. In Sec. V.1, we propose several properties for $H_{E}$, $H_{S\!E}$ and $\rho_{E}(t_{I})$ that guarantee the commutation relation between a parity kick $K_{P}$ and $\bar{T}(t_{b},t_{a})$, which is proved with the parities of the total excitation numbers in the ket and bra vectors of the environment initial states. However, general conditions that guarantee $K_{j}\bar{T}(t_{b},t_{a})=\bar{T}(t_{b},t_{a})K_{j}$ for an arbitrary dynamical decoupling process are still open.